Find the solution of


(D^2-4D+4)y=12x+2e^(3x)

1
Expert's answer
2022-02-07T15:16:10-0500

The given DE is "\displaystyle
(D^2-4D+4)y=12x+2e^{3x}".

Now, the auxiliary equation is;

"m^2-4m+4=0\Rightarrow(m-2)^2=0\Rightarrow m=2" (twice)

"\therefore y_c=(c_1+c_2x)e^{2x}", where "c_1,c_2" are constants.


Next, to obtain the particular integral;

"y_p=(D^2-4D+4)^{-1}12x+\frac{1}{D^2-4D+4}2e^{3x}\\
\quad\ =3\left[1-D+\frac{D^2}{4}\right]^{-1}x+\frac{1}{9-12+4}2e^{3x}\\
\quad\ =3\left[1+D-\frac{D^2}{4}\right]x+2e^{3x}\\
\quad\ =3\left[x+D(x)-\frac{D^2}{4}(x)\right]+2e^{3x}\\
\quad\ =3[x+1-0]+2e^{3x}=3x+3+2e^{3x}"


Thus, the general solution is;

"y=y_c+y_p=(c_1+c_2x)e^{2x}+3x+3+2e^{3x}"